If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days? Can you show me how to work it out?
Answer by babynarah2007(12)

If 6 men can do a piece of work in 14 days, how many men are needed to do the work in 21 days? Can you show me how to work it out?
Factors 21 and 14 have greatest common factor (GCF) of 7. Reducing fraction.
As we know; 6 men "X" work at 14 days
9 men "x" work must be 21 days
the expression: 6 men 'X' work at 14 days
-----? 'x' work at 21 days
so that, we need to solve by how many men?
equation: 21*6/14
by Factors 21 and 14 have greatest common factor (GCF) of 7. Reducing fraction.
= 21*6/14
= 3*6/2
= 3*3/(1)note we can take out '1'
= 3*3
= 9 men
done! good luck!!
by narah